# STORYANCHORS: Generating Consistent Multi-Scene Story Frames for Long-Form Narratives

### v1

@@ 1 Introduction
An alternative and more effective approach is story-frame generation, where story frames are first created using methods similar to text-to-image techniques, and then image-to-video models are used to interpolate intermediate frames to construct the full video. This strategy offers a more controlled and flexible way to generate story-driven content, allowing for better narrative progression across multiple scenes. However, existing story-frame generation methods face significant limitations. IC-LoRA[6], for example, is highly constrained by its ability to generate only limited frames at a time, severely limiting its flexibility and ultimately degrading the quality of the story frames. Meanwhile, methods like StoryDiffusion[7], VideoGen-of-Thought[8], and One Prompt One Story[9], while more flexible, heavily focus on character-centric shots, leading to repetitive perspectives and weak story development. Additionally, as the number of story frames increases, challenges like consistency, semantic alignment, and narrative coherence become more pronounced, further diminishing the overall video quality. Recent multimodal large language models (MLLMs) like GPT-4o[10] can generate multiple story frames but require detailed prompts and often struggle with consistency and producing the requested number of frames. Over time, reliance on historical context leads to accumulating inconsistencies and a lack of global narrative coherence, hindering the creation of high-quality, multi-scene story-driven frames.
@@ 2 Approach
In the following sections, we introduce the core components of our StoryAnchors framework, including Bidirectional Story Frame Prediction, Multi-Event Story Frame Labeling, and Progressive Story Frame Training. These components work together to enable the generation of coherent, story-driven videos characterized by high narrative richness, multi-scene continuity, and consistent thematic progression, with flexible scene transitions and strong temporal coherence across diverse story elements.
@@ 2.1 Bidirectional Story Frame Prediction
Our framework is built upon the pre-trained Step-Video-T2V-4B model, with modifications made in both the input stage and the AdaLN module (as illustrated in Figure 2). The original input latent shape for T2V is $[f,c,h,w]$, where $f$, $c$, $h$, and $w$ represent the number of frames, channels, height, and width. In our Story Frame Prediction Network, we modify this by including condition frames or zero padding as conditional hidden states. Similar to TI2V models [12, 13], we expand the input channels to accommodate additional channels for the condition frames. Moreover, a mask channel with a shape of $[f,1,h,w]$ is introduced, marking the frames to be generated versus the condition frames. A zero value in the mask indicates that the frame should be generated, while other values correspond to condition frames. This results in a concatenated input latent of shape $[f,2c+1,h,w]$, and we modify the patch embedding module of the DiT architecture to handle this expanded input structure.
In StoryAnchors, we adopt a flexible framework that allows users to influence the story frame generation by incorporating diverse condition embeddings. As shown in Figure 2, these embeddings include:
@@
  • Frames Motion to adjust the dynamic degree of the video,
  • Frames Resolution to define the input video’s width and height, aiding the model in understanding the original resolution and supporting mixed-resolution training,
@@
These embeddings are added to the Timestep Embedding before being injected into the AdaLN modules to influence the stroy frame generation process.
After generating the story frames, StoryAnchors employs a text driven image-to-video interpolation model, such as Step-Video-TI2V, to synthesize intermediate frames between any two consecutive story frames. This enables smooth transitions between scenes and the generation of multi-shot videos with strong temporal consistency and rich narrative progression. By simply concatenating the generated video chunks, we can create a long-form video that retains both visual coherence and story-driven structure.
@@ 2.3 Progressive Story Frame Training
Stage 1: Pretraining with Global Caption Due to the inherent difficulty and cost of fine-grained multi-event annotation, we initiate training using an automatically annotated dataset containing 2 million frame-caption pairs. These captions are generated by a Step-LLM model with global prompts, which capture high-level event transitions across entire videos without explicit segmentation into discrete prompts. This initial step allows the model to adapt from a frame-based representation to a story frame generation process. Although the captions lack fine temporal granularity, they maintain the overall event chronology, which is essential for developing the model’s understanding of narrative flow.
@@
Stage 3: Story Anchors Prediction for Story Continuation To enhance the model’s ability to generate temporally extended captions and support story continuation, we introduce a keyframe prediction module inspired by the TI2V paradigm. Specifically, we train the model to predict future story frames conditioned on the first several frames and their associated prompts, allowing the system to generate plausible event progressions from minimal visual input. This stage strengthens the model’s ability to generate more detailed, extended narratives, supporting the generation of complex anchor sequences over longer horizons.
3 Related Work
Multimodal Large Language Models.
While Multimodal Language Models (MLLMs) [15, 16, 17] were initially developed for multimodal understanding, recent research has increasingly demonstrated the potential of MLLMs in video comprehension [18, 19]. More recently, MLLMs have begun to expand toward generative tasks, including image synthesis [20, 21], and even frame-by-frame keyframe generation [7] for constructing videos. Notably, models like GPT-4o [10] and PikaFrames [22] have gained attention for their ability to generate videos one frame at a time, treating each frame as a high-quality, text-aligned keyframe. These models demonstrate impressive aesthetic fidelity and short-term consistency, making them effective for localized content generation. However, due to their autoregressive nature, they typically rely only on past visual context, limiting their ability to perform global planning across entire video sequences. As a result, such approaches often struggle with scene diversity, temporal coherence, and narrative structure, frequently producing short, close-up shots rather than long-form, multi-shot stories.
Scene-level Video Generation.
Recent advances in diffusion-based models[23], such as SoRA[24], have significantly boosted video generation through 3D modeling and full-attention mechanisms, inspiring both commercial (e.g., Kling[25], Gen-3[26]) and open-source systems (e.g., HunyuanVideo[27], Wan2.1[13], StepVideo[11, 12]). Current scene-level generation approaches mainly fall into two categories: keyframe-based and autoregressive methods.
On the other hand, autoregressive approaches such as FreeNoise[3], StreamingT2V[2], and CasusVid[5] extend video length by generating frames sequentially, often leveraging rescheduled noise or window-based temporal attention. More recent efforts like MinT[31] and DFoT[4] enhance consistency with time-conditioned prompts or history-guided rollout. Nonetheless, these methods still lack explicit global planning of key visual anchors and struggle to maintain coherent narratives across scenes.
To address these challenges, we propose a story frame centric framework that emphasizes globally consistent, semantically dense keyframes as the foundation for generating coherent, multi-scene, minute-scale videos.
@@ 4.1 Implement Details
We adopt the Step-Video-T2V-4B variant [11, 32] as our base model, with its text encoder built upon Step-LLM [11]. For the VAE component, we employ WanVAE [13]. During the text annotation process, keyframes were uniformly extracted from the original videos at intervals of 69 or 139 frames. An initial subset of the data was annotated using the Gemini system. These annotations were then used to bootstrap our in-house vision-language large model, which was subsequently employed to scale up the annotation process.
@@ 4.2 Comparisons of Consistent Story Frames Generation
Qualitative Comparisons. Figure 4 presents a qualitative comparison between our method (StoryAnchors) and several baselines for consistent story frame generation, including StoryDiffusion [7], VideoGen-of-Thought [8], One Prompt One Story [9], and GPT-4o. Each sequence illustrates how a model interprets a given prompt to construct a coherent visual narrative. Among the methods evaluated, our model, along with VideoGen-of-Thought and One Prompt One Story, consistently produces sequences with higher thematic and background coherence. Notably, StoryAnchors demonstrates greater shot diversity, featuring a broader range of camera angles that enhance visual storytelling. In contrast, GPT-4o often defaults to mid-shots, resulting in limited compositional variety and reduced narrative engagement. While StoryDiffusion generates relatively diverse shots, it struggles with maintaining logical event progression and temporal consistency, particularly in complex scenes. In comparison, our model exhibits stronger narrative coherence and a clearer depiction of event dynamics and emotional transitions. Moreover, GPT-4o frequently suffers from detail inconsistencies and color shifts, especially as the number of generated keyframes or dialogue turns increases. Our model, by contrast, maintains both visual fidelity and consistency across frames, while delivering comparable or superior performance in plot development and scene richness.
@@
Table 1: Comparison results with story frames generated by GPT-4o.
The quantitative results are in line with our qualitative findings. When generating 10 frames, StoryAnchors achieves strong overall performance, particularly excelling in subject consistency and scene diversity. For other open-source methods, increasing the number of generated keyframes tends to enhance narrative richness and variation. However, this improvement often comes at the cost of reduced consistency, as these models struggle to maintain coherent subject representations and stable scene layouts beyond the 10-frame mark. We also conduct a comparative evaluation against GPT-4o. As shown in Table 1, GPT-4o meets the resolution and instruction-following requirements in 71.4% of cases. While StoryAnchors yields slightly lower scores in NQ and PFA, it delivers comparable overall results, with several metrics showing parity between the two models.
4.3 Story Frames Editing and Temporal Extension
@@ 4.4 Effectiveness of Multi-Event Prompt
The Global Prompt method treats the video as a single, continuous sequence and generates frames based on a high-level description without distinguishing individual events. This approach results in significant challenges in maintaining event completeness and temporal order. Without explicit event segmentation, the model struggles to generate frames that are consistent with the sequence of events, leading to fragmented and disordered narratives. In contrast, our Multi-Event Description approach segments the video into distinct events, with each event described in detail. This method ensures better narrative consistency, with clear transitions between events and better alignment between the prompt and generated frames.
@@ 5 Conclusion
In this work, we introduce STORYANCHORS, a novel framework for generating consistent, multi-scene videos by predicting a sequence of globally coherent story frames. By leveraging bidirectional context modeling, STORYANCHORS captures long-range dependencies such as recurring characters and evolving scenes, enabling flexible, controllable narrative design. The framework offers a powerful solution for long-form video generation by decoupling the keyframe prediction from intermediate frame synthesis, allowing for scalable video creation with strong temporal consistency. Beyond video synthesis, STORYANCHORS highlights the potential of keyframe-based methods to promote both video generation and understanding. By using story frames as high-level anchors, we believe this approach can advance the integration of generation and comprehension, paving the way for more coherent and expressive video content. Future work will focus on further improving the flexibility of keyframe generation and exploring the full potential of this integrated framework in various storytelling applications.
@@ Appendix A Evaluation Metrics
  • A: Narrative Quality (NQ): Evaluates how clearly and coherently the generated frames convey a complete story arc.
@@
2 I provide the second set of images from the same video.
@@
3 [Label Prompt] [Assistant’s Response based on Label Prompt]
Narrative Quality (NQ), Subject Consistency (SC), Prompt-Frame Alignment (PFA), Shot Diversity (ShotD), and Scene Diversity (SceneD)

### v2

@@ 1 Introduction
An alternative and more effective approach is story-frame generation, where story frames are first created using methods similar to text-to-image techniques, and then image-to-video models are used to interpolate intermediate frames to construct the full video. This strategy offers a more controlled and flexible way to generate story-driven content, allowing for better narrative progression across multiple scenes. However, existing story-frame generation methods face significant limitations. IC-LoRA[6], for example, is highly constrained by its ability to generate only limited frames at a time, severely limiting its flexibility and ultimately degrading the quality of the story frames. Meanwhile, methods like StoryDiffusion[7], VideoGen-of-Thought[8], and One Prompt One Story[9], while more flexible, heavily focus on character-centric shots, leading to repetitive perspectives and weak story development. Additionally, as the number of story frames increases, challenges like consistency, semantic alignment, and narrative coherence become more pronounced, further diminishing the overall video quality.
Recent multimodal large language models (MLLMs) like GPT-4o[10] can generate multiple story frames but require detailed prompts and often struggle with consistency and producing the requested number of frames. Over time, reliance on historical context leads to accumulating inconsistencies and a lack of global narrative coherence, hindering the creation of high-quality, multi-scene story-driven frames.
@@
2 Related Work
Multimodal Large Language Models.
While Multimodal Language Models (MLLMs) [13, 14, 15] are initially developed for multimodal understanding, recent researches increasingly demonstrate the potential of MLLMs in video comprehension [16, 17]. More recently, MLLMs have begin to expand toward generative tasks, including image synthesis [18, 19], and even frame-by-frame keyframe generation [7] for constructing videos. Notably, models like GPT-4o [10] and PikaFrames [20] have gain attention for their ability to generate videos one frame at a time, treating each frame as a high-quality, text-aligned keyframe. These models demonstrate impressive aesthetic fidelity and short-term consistency, making them effective for localized content generation. However, due to their autoregressive nature, they typically rely only on past visual context, limiting their ability to perform global planning across entire video sequences. As a result, such approaches often struggle with scene diversity, temporal coherence, and narrative structure, frequently producing short, close-up shots rather than long-form, multi-shot stories.
Scene-level Video Generation.
Recent advances in diffusion-based models[21], such as SoRA[22], have significantly boosted video generation through 3D modeling and full-attention mechanisms, inspiring both commercial (e.g., Kling[23], Gen-3[24]) and open-source systems (e.g., HunyuanVideo[25], Wan2.1[26], StepVideo[11, 12]). Current scene-level generation approaches fall mainly into two categories: keyframe-based and autoregressive methods.
On the other hand, autoregressive approaches such as FreeNoise[3], StreamingT2V[2], and CasusVid[5] extend video length by generating frames sequentially, leveraging rescheduled noise or temporal attention. More recent efforts like MinT[30] and DFoT[4] enhance consistency with time-conditioned prompts or history-guided rollout. Nonetheless, these methods still lack explicit global planning of key visual anchors and struggle to maintain coherent narratives across scenes.
To address these challenges, we propose a story frame centric framework that emphasizes globally consistent and semantically rich frames as the structural backbone for generating coherent, multi-scene, minute-scale videos. These story frames serve as high-level anchors that guide both visual content and narrative flow, enabling precise control over scene transitions, and temporal progression, and thematic continuity throughout the video.
@@ 3 Approach
In following sections, we introduce the core components of our StoryAnchors framework, including Bidirectional Story Frame Prediction, Multi-Event Story Frame Labeling, and Progressive Story Frame Training. These components enable the generation of coherent, story-driven videos characterized by high narrative richness, multi-scene continuity, and consistent thematic progression, with flexible scene transitions and strong temporal coherence across diverse story elements.
@@ 3.1 Bidirectional Story Frame Prediction
Our framework is built upon the pre-trained Step-Video-T2V-4B model, with modifications made in both the input stage and the AdaLN module (as illustrated in Figure 2). The original input latent shape for T2V is $[f,c,h,w]$, where $f$, $c$, $h$, and $w$ represent the number of frames, channels, height, and width. In our Story Frame Prediction Network, we modify this by including condition frames or zero padding as conditional hidden states. Similar to TI2V models [12, 26], we expand the input channels to accommodate additional channels for the condition frames. Moreover, a mask channel with a shape of $[f,1,h,w]$ is introduced, marking the frames to be generated versus the condition frames. A zero value in the mask indicates that the frame should be generated, while other values correspond to condition frames. This results in a concatenated input latent of shape $[f,2c+1,h,w]$, and we modify the patch embedding module of the DiT architecture to handle expanded input structure.
StoryAnchorsadopt a flexible framework that allows users to influence the story frame generation by incorporating diverse condition embeddings. Shown in Figure 2, these embeddings include:
@@
  • Frames Motion to adjust the dynamic degree of the video.
  • Frames Resolution to define the input video’s width and height, aiding the model in understanding the original resolution and supporting mixed-resolution training.
@@
These embeddings are added to the Timestep Embedding before being injected into the AdaLN modules to influence the story frame generation process.
After generating the story frames, StoryAnchors employs a text driven image-to-video interpolation model, such as Step-Video-TI2V, to synthesize intermediate frames between any two consecutive story frames. This enables smooth transitions between scenes and the generation of multi-shot videos with strong temporal consistency and rich narrative progression.
@@ 3.3 Progressive Story Frame Training
Stage 1: Pretraining with Global Caption Due to the inherent difficulty and cost of fine-grained multi-event annotation, we initiate training using an automatically annotated dataset containing 2 million frame-caption pairs. These captions are generated by a Step-LLM model with global prompts, which capture high-level event transitions across entire videos without explicit segmentation into discrete prompts. This initial step allows the model to adapt from a frame-based representation to a story frame generation process. Although the captions lack fine temporal granularity, they maintain the overall event chronology, which is essential for developing the model’s understanding of narrative.
@@
Stage 3: Story Anchors Prediction for Story Expansion To enhance the model’s ability to generate temporally extended captions and support story continuation, we introduce a keyframe prediction module inspired by the TI2V paradigm. Specifically, we train the model to predict future story frames conditioned on the first several frames and their associated prompts, allowing the system to generate plausible event progressions from minimal visual input. This stage strengthens the model’s ability to generate more detailed, supporting the generation of complex anchor sequences over longer horizons.
@@ 4.1 Implement Details
We adopt the Step-Video-T2V-4B variant [11, 32] as our base model, with its text encoder built upon Step-LLM [11]. For the VAE component, we employ WanVAE [26]. During the annotation, keyframes were uniformly extracted from the original videos at intervals of 69 or 139 frames. An initial subset of the data was annotated using the Gemini system. Each video is center-cropped and resized to 480×768. The Adam optimizer with a learning rate of 1e-5 is used for each training stage. A total batch size of 256 is distributed across 64 H100 GPUs.
@@ 4.2 Comparisons of Consistent Story Frames Generation
Qualitative Comparisons. Figure 4 presents a qualitative comparison between our method (StoryAnchors) and several baselines for consistent story frame generation, including StoryDiffusion [7], VideoGen-of-Thought [8], One Prompt One Story [9], and GPT-4o. Each sequence illustrates how a model interprets a given prompt to construct a coherent visual narrative. Among the methods evaluated, our model, along with VideoGen-of-Thought and One Prompt One Story, consistently produces sequences with higher thematic and background coherence. Notably, StoryAnchors demonstrates greater shot diversity, featuring a broader range of camera angles that enhance visual storytelling. In contrast, GPT-4o often defaults to mid-shots, resulting in limited compositional variety and reduced narrative engagement. While StoryDiffusion generates relatively diverse shots, it struggles with maintaining logical event progression and temporal consistency, particularly in complex scenes. In comparison, our model exhibits stronger narrative coherence and a clearer depiction of event dynamics and emotional transitions. Moreover, GPT-4o frequently suffers from detail inconsistencies and color shifts, especially as the number of generated keyframes or dialogue turns increases. Our model, by contrast, maintains both visual fidelity and consistency across frames, while delivering comparable or superior performance in narrative quality and scene richness.
Table 1: Comparison results (Win/Tie/Loss) based on story frames generated by GPT-4o.
@@
The quantitative results are in line with our qualitative findings. When generating 10 frames, StoryAnchors achieves strong overall performance, particularly excelling in subject consistency and scene diversity. For other open-source methods, increasing the number of generated keyframes tends to enhance narrative richness and variation. However, this improvement often comes at the cost of reduced consistency, as these models struggle to maintain coherent subject representations and stable scene layouts beyond the 10-frame mark. We also conduct a comparative evaluation against GPT-4o. As shown in Table LABEL:tab:gpt4o_single_turn_transposed, GPT-4o meets the resolution and instruction-following requirements in 71.4% of cases. While StoryAnchors yields slightly lower scores in NQ and PFA, it delivers comparable overall results, with several metrics showing parity between the two models.
@@ 4.4 Effectiveness of Multi-Event Prompt
Figure 12: Comparison between global prompt and multi-event descriptions. Each subfigure presents a generated video sequence using (top) a single global prompt and (bottom) multi-event prompt.
The Global Prompt method treats the video as a single, continuous sequence and generates frames based on a high-level description without distinguishing individual events. This approach results in significant challenges in maintaining both event completeness and temporal order. For instance, in Figure LABEL:fig:prompt_comparison(a), the global-prompt-based generation omits important moments such as the squirrel jumping over the stone and the girl’s walk along the lakeside, both of which are crucial for a complete depiction of the narrative. Similarly, in Figure LABEL:fig:prompt_comparison(b), the global prompt introduces large numbers of butterflies too early and misplaces the butterfly-shaped map at the end, despite it being a key discovery that initiates the journey, thereby disrupting the logical event progression.
In contrast, our Multi-Event Description approach explicitly segments the story into structured events, each guided by detailed descriptions. This enables more accurate and coherent visual storytelling, ensuring narrative consistency, event completeness, and temporally aligned generation. As shown in both examples, the multi-event method captures all significant moments in the correct sequence, resulting in a more faithful and interpretable depiction of the story.
@@ 5 Conclusion
In this work, we introduce STORYANCHORS, a novel framework for generating consistent, multi-scene videos by predicting a sequence of globally coherent story frames. By leveraging bidirectional context modeling, STORYANCHORS captures long-range dependencies such as recurring characters and evolving scenes, enabling flexible, controllable narrative design. The framework offers a powerful solution for long-form video generation by decoupling the keyframe prediction from intermediate frame synthesis, allowing for scalable video creation with strong temporal consistency. Beyond video synthesis, STORYANCHORS highlights the potential of keyframe-based methods to promote both video generation and understanding. By using story frames as high-level anchors, we believe this approach can advance the integration of generation and comprehension, paving the way for more coherent and expressive video content.
Limitation and broader impact. While StoryAnchors demonstrates strong capabilities in generating semantically consistent and temporally coherent story frames, its current reliance on TI2V-based generation. This limits the model’s ability to adapt to complex scenes or tasks, often resulting in weak transitions and inconsistent narrative continuity. Future work will explore more effective global modeling of all anchors to enhance temporal coherence across extended video sequences. Furthermore, we must use the method responsibly to prevent any negative social impacts, such as the creation of misleading fake videos.
@@ Appendix A Evaluation Metrics
  • Narrative Quality (NQ): Evaluates how clearly and coherently the generated frames convey a complete story arc.
@@
2 I provide the second set of story frames from the same video.
@@
3 [Evaluation Criteria] [Assistant’s Response based on Evaluation Criteria]
Appendix B Evaluation Criteria
We present the evaluation criteria for all metrics, including Narrative Quality (NQ), Subject Consistency (SC), Prompt-Frame Alignment (PFA), Shot Diversity (ShotD), and Scene Diversity (SceneD).